OpenStudy (anonymous):

Find all zeros 5x^4+126x^2+25 Need help setting up.

OpenStudy (anonymous):

Make the substitution y=x^2 Doing this you get: \[5x^4+126x^2+25=5y^2+126y+25\] Now factor this: \[5y^2+126y+25=(5y+1)(y+25)\] But remember, y=x^2, so: \[(5x^2+1)(x^2+25)\] And this doesn't factor any further on the real numbers.

OpenStudy (anonymous):

So there are no x-intercepts. (zeros)

OpenStudy (mertsj):

The fact that there are no zeros can be discovered by looking at the equation. Since x^4 has a minimum value of 0 and x^2 has a minimum value of 0, the y value will never be less than 25 so it could never cross the x axis. Thus no zeros.

OpenStudy (anonymous):

Thankk youu!

OpenStudy (mertsj):

yw